OpenStudy (anonymous):

Express the ratio in simplest form. 18 : 12

OpenStudy (kidrah69):

\[\frac{ 18 }{ 12 }\]

OpenStudy (kidrah69):

simplyfy it :)

OpenStudy (kendricklamar2014):

^

OpenStudy (kendricklamar2014):

Find the least common multiples (LCM) of both 18 and 12 and divide them to their lowest number.

OpenStudy (kendricklamar2014):

LCM of 18 and 12 is 6 so do:\[\frac{ 18 }{ 6 } = ?\]\[\frac{ 12 }{ 6 } = ?\]